A gas cylinder can safely handle pressures up to 75.0 atm. If the gas pressure in the cylinder is 63.3 atm at 15 ◦C, what is the highest safe temperature for this cylinder? Assume ideal-gas behavior and V, n = constant. T(K) = T(◦C) + 273.

(A) 17.8 ◦C (B) 68 ◦C (C) 243 ◦C (D) 291 ◦C (E) 341 ◦C
Answer #1

Given:

Ti = 15.0 oC

= (15.0+273) K

= 288 K

Pi = 63.3 atm

Pf = 75.0 atm

use:

Pi/Ti = Pf/Tf

63.3 atm / 288.0 K = 75.0 atm / Tf

Tf = 341 K

= (341 - 273) oC

= 68 oC

Answer: B